Using the Runge-Kutta fourth-order method, obtain a solution to

dx/dt=f(t,x,y)=xy^3+t^2; dy/dt=g(t,x,y)=ty+x^3

for t= 0 to t= 1 second. The initial conditions are given as x(0)=0, y(0) =1. Use a time increment of 0.2 seconds. Do hand calculations for t = 0.2 sec only.
Answer #1

Cr Cǐven differential equation System Ys ot h=o-2 and The 4t order Rune-kuta fermulals for q system of 2 oDEre Lo Where 2- Кр= (M) (0.01) = o gリ Кр %リ 1.0 -6.2) f (0-1, o-ml ^= (ot) (。.loIn) ะ ‘.e2。ユ 0 f (6-2 , 0.00 2206 , IRAJ う 3 6 2) Ce.go4c49 =o.oYo808 3 Henceand Hence,Solution at t-o . 2